OpenStudy (anonymous):

Two triangles can be formed with the given information. Use the Law of Sines to solve the triangles. A = 55°, a = 12, b = 14

OpenStudy (anonymous):

@agent0smith

OpenStudy (agent0smith):

Try drawing it first

OpenStudy (anonymous):



OpenStudy (agent0smith):



OpenStudy (anonymous):

I HAVE B = 72.9°

OpenStudy (agent0smith):

SinB/14 = sin55/12 so B = 72.9 OR 180-72.9

OpenStudy (anonymous):

yes i got that

OpenStudy (anonymous):

and for c = 12.5;

OpenStudy (agent0smith):

C = 52.1 or 17.9 then c/sinC = 12/sin55 Solve for c using both above angles

OpenStudy (anonymous):

c = 11.6 or c = 4.5

OpenStudy (anonymous):

Im i doing it wrong?

OpenStudy (agent0smith):

idk, i will be back later c/sin52.1 = 12/sin55 c/sin17.9 = 12/sin55 solve them and you're done

OpenStudy (anonymous):

@agent0smith

OpenStudy (anonymous):

@e.mccormick

OpenStudy (agent0smith):

It's all in the last post.